# Systematic discrepancies between reference methods for non-covalent interactions within the S66 dataset

### v1

@@ Abstract
The accurate treatment of non-covalent interactions is necessary to model a wide range of applications, from molecular crystals to surface catalysts to aqueous solutions and many more. Quantum diffusion Monte Carlo (DMC) and coupled cluster theory with single, double and perturbative triple excitations [CCSD(T)] are considered two widely-trusted methods for treating non-covalent interactions. However, while they have been well-validated for small molecules, recent work has indicated that these two methods can disagree by more than 7.57.57.5\,7.5kcal/mol for larger systems. The origin of this discrepancy remains unknown. Moreover, the lack of systematic comparisons, particularly for medium-sized complexes, has made it difficult to identify which systems may be prone to such disagreements and the potential scale of these differences. In this work, we leverage the latest developments in DMC to compute interaction energies for the entire S66 dataset, containing 66 medium-sized complexes with a balanced representation of dispersion and electrostatic interactions. Comparison to previous CCSD(T) references reveals systematic trends, with DMC predicting stronger binding than CCSD(T) for electrostatic-dominated systems, while the binding becomes weaker for dispersion-dominated systems. We show that the relative strength of this discrepancy is correlated to the ratio of electrostatic and dispersion interactions, as obtained from energy decomposition analysis methods. Finally, we pinpoint systems in the S66 dataset where these discrepancies are particularly prominent, offering cost-effective benchmarks to guide future developments in DMC, CCSD(T) as well as the wider electronic structure theory community.
@@ I Introduction
The origin of the discrepancy between DMC and CCSD(T) for large dispersion-bound molecules is a topic of current debate, Schäfer et al. (2024); Fishman et al. (2024); Lambie et al. (2024); Lao (2024) particularly on the validity of the perturbative triples (T) contribution in CCSD(T). Schäfer et al. Schäfer et al. (2024) have suggested that part of this discrepancy arises from missing contributions in (T) that can be accounted by the (cT) approach. In addition, Semidalas et al. Semidalas et al. (2022) have reported non-trivial discrepancies between CCSD(T) and post-CCSD(T) methods such as CCSDT(Q). Conversely, Lambie et al. Lambie et al. (2024) have found that CCSD(T) does not differ significantly against CCSDT(Q) using the Pariser-Parr-Pople (PPP) model Pariser and Parr (1953a, b); Pople (1953) for large conjugated systems. Similarly, Fishman et al. Fishman et al. (2024) and Lao Lao (2024) report only a slight overbinding of CCSD(T) against CCSDT(Q) that cannot explain the discrepancy against DMC.
@@ II.1 Diffusion Monte Carlo
A detailed description of the DMC method can be found in Ref. 9. In this work, we compute fixed-node DMC interaction energies by using the CASINO codeNeeds et al. (2020). We use energy-consistent correlated electron pseudopotentialsTrail and Needs (2017) (eCEPP) with the determinant locality approximation (DLA)Zen et al. (2019). The trial wave-functions were of the Slater–Jastrow type with single Slater determinants, and the single-particle orbitals obtained from DFT local-density approximation (LDA) plane-wave calculations performed with PWscfGiannozzi et al. (2009, 2020) using an energy cut-off of 600 Ry and re-expanded in terms of B-splinesAlfè and Gillan (2004). The Jastrow factor included a two-body electron–electron (e–e) term, two-body electron–nucleus (e–n) terms, and three-body electron–electron–nucleus (e–e–n) terms. The variational parameters of the Jastrow have been optimized by minimizing the variance of each system. The final DMC estimates of $\Delta E_{\textrm{int.}}$ were extrapolated towards the zero time step limit ($\tau\to 0$) by making a cubic fit to a series of time step estimates from 0.110.110.11\,0.11au down to 0.0030.0030.003\,0.003au. We have estimated errors on this fit with two approaches: either by taking the $1\sigma$ standard deviation of the fit at zero time step or the difference w.r.t. the zero time step limit given by a linear fit from the two smallest (0.010.010.01\,0.01au and 0.0030.0030.003\,0.003au) time steps; the latter accounts for systems where the time step behavior may change close to the zero time step limit. The final error estimate is taken to be the larger of the two, as discussed in Sec. S2.4 of the supplementary material.
@@ II.2 Coupled Cluster Theory
Schäfer et al. Schäfer et al. (2024) have recently demonstrated that there exists an empirical relationship between the (cT) and the (T) correlation contributions to the total energy using the CCSD and MP2 correlation energies. The resulting (cT)-fit is of the form:
@@ III Results and Discussion
This dataset comprises of a diverse range of interactions and we have separated the systems according to the original S66 categories of hydrogen-bonded, dispersion-bonded and “mixed”-character systems in Figs. 12 and 3, respectively. The corresponding dimer complexes are visualized in Fig. S1 of the supplementary material. We report the DMC estimate of the interaction energy $\Delta E_{\textrm{int.}}$ above the label of each S66 complex. In all cases, the rrors on $\Delta E_{\textrm{int.}}$ estimates are below 0.260.260.26\,0.26kcal/mol, with the majority below 0.100.100.10\,0.10kcal/mol, facilitating reliable comparisons to other CCSD(T).
@@
The strength of $\Delta E_{\textrm{int.}}$ varies significantly across the systems, from as large as $-20.17\pm 0.11\,$kcal/mol for complex 20 (acetic acid dimer) to as weak as $-1.11\pm 0.06\,$kcal/mol for complex 30 (benzene-ethene dimer), being stronger in the H-bonded systems. With gray crosses, we plot the difference between DMC and CCSD(T) estimates (as described in the Methods) for the three classes of interactions. We use DMC as the reference (i.e., zero), and plot blue vertical bars along the horizontal zero-axis representing the errors on the DMC estimates. There is overall excellent agreement, with a mean absolute deviation (MAD) of 0.210.210.21\,0.21kcal/mol across the entire S66 dataset. We find systematic trends in the differences between CCSD(T) and DMC, with CCSD(T) predicting weaker binding compared to DMC for hydrogen-bonded systems in Fig. 1, with an MAD of $\sim 0.24\,$kcal/mol, while predicting a stronger binding for dispersion dominated systems in Fig. 2, with an MAD of $\sim 0.24\,$kcal/mol. For the “mixed” character systems in Fig. 3, the MAD is lower at 0.140.140.14\,0.14kcal/mol.
Figure 2: Comparison between DMC interaction energies $\Delta E_{\textrm{int.}}$ calculated in the present work against CCSD(T) for a subset of systems in the S66 dataset dominated by dispersion interactions. The first 6 dimers are $\pi{-}\pi$ stacked. Refer to the caption of Fig. 1 for the plot details. The CCSD(T) estimate is taken as an average from three previous calculations Řezáč, Riley, and Hobza (2011b); Kesharwani et al. (2018); Nagy et al. (2023), with corresponding standard deviation as error. Additional CCSD(cT)-fit estimates are are reported with golden crosses. These are calculated by scaling the CCSD(T) estimates based on their MP2 and CCSD contributions with the approach described in Ref. 50.
@@
The stronger binding of DMC over CCSD(T) has not been (systematically) reported before, with the acetic acid dimer (ID 20) giving the maximum deviation of 0.80.80.8\,0.8kcal/mol across all S66 systems. Within Sec. S7 of the supplementary material, we have confirmed that the computed $\Delta E_{\textrm{int.}}$ estimate ($-20.17\pm 0.11\,$kcal/mol) does not depend on the chosen pseudopotential (eCEPP), localization scheme (DLA) or trial wave-function (LDA). For example, we have also performed all-electron calculations, giving an estimate of $-20.19\pm 0.11\,$kcal/mol (at $\tau{=}0.0015\,$au that is within the statistical uncertainties of our original estimate. We have also performed tests using PBE and PBE0 trial wave-functions, showing that the DMC $\Delta E_{\textrm{int.}}$ has a negligible (${<}0.15\,$kcal/mol) dependence on the nodal surface for the DFAs considered. Furthermore, we also computed estimates for two-other localization schemes: T-move and determinant localization T-move, both of which were within the statistical uncertainties of our original estimate. It should be noted that while the absolute value of the difference can be significant for some hydrogen-bonded systems, the relative difference (normalized against the DMC $\Delta E_{\textrm{int.}}$) is significantly smaller, with a mean relative difference of 2.45% compared to 8.17% for the dispersion-dominated systems (see Sec. S6 of the supplementary material).
@@
Figure 4: Error decomposition analysis. We report the absolute difference between DMC and CCSD(T) relative to the DMC values, i.e. $\left(E_{\mathrm{DMC}}-E_{\mathrm{CCSD(T)}}\right)/\left|E_{\mathrm{DMC}}\right|$, as a function of the natural logarithm of the electrostatic (ELST) to dispersion (DISP) ratio contribution to the binding energy. The ELST to DISP ratio is determined from the SAPT analysis from Ref. 93. The color code is red for H-bonded systems (ID from 1 to 23), blue for dispersion dominated systems (ID from 24 to 46), and green for mixed systems (ID from 47 to 66).
@@
The weaker binding of DMC over CCSD(T) for dispersion-dominated systems is now relatively well-documented Al-Hamdani et al. (2021); Schäfer et al. (2024), and there is evidence that it can be improved by replacing the perturbative triples (T) contribution with the recent (cT) Masios et al. (2023) contribution. We plot the difference between an empirically CCSD(cT)-fit formulation as yellow crosses for the dispersion-dominated systems. In all cases, CCSD(cT) has a weaker binding than CCSD(T), leaning closer towards DMC, leading to an MAD of 0.090.090.09\,0.09kcal/mol and mean relative difference of 3.43% However, this does not fully resolve the discrepancies across all of the dispersion systems, with significant discrepancy of ${\sim}0.4\,$kcal/mol remaining for the uracil-cyclopentane dimer (ID 42). Such a significant discrepancy makes this a worthwhile system to investigate further and could give clues on remaining discrepancies between DMC and CCSD(T) observed in Ref. 50. The analysis reported above highlights an important outcome of this work: the identification of smaller, simpler systems that show notable discrepancies between DMC and coupled cluster methods. Specifically, we found a discrepancy of approximately ${\sim}0.9\,$kcal/mol for the hydrogen-bonded acetic acid dimer (ID 20) and ${\sim}0.5\,$kcal/mol for the dispersion-dominated uracil–cyclopentane dimer (ID 43), which contain 64 and 98 (total) electrons, respectively. These medium-sized systems represent an almost tenfold reduction in electron count compared to the larger $\mathrm{C}_{60}@[6]\mathrm{CPPA}$ buckyball-ring system (672 electrons) studied in Ref. 49. Thus, they might offer practical, cost-effective models for further exploring the discrepancy between DMC and coupled-cluster.
@@
Finally, we now focus on the difference between DMC and CCSD(T) as a function of the dispersion and electrostatic contribution to the interaction energy. In particular, we find that the relative differences between DMC and CCSD(T) for each system within the S66 dataset can be correlated to the relative strength of the dispersion and electrostatic interactions that make up its $\Delta E_{\textrm{int.}}$. We used the Symmetry Adapted Perturbation Theory (SAPT) calculations from Burns et al.Burns et al. (2017), which decomposes $\Delta E_{\textrm{int.}}$ into contributions from electrostatics (ELST), exchange, induction, and dispersion (DISP). Notably, we show in Fig. 2 that there is a strong linear trend ($R^{2}{=}0.78$) between the natural logarithm of the ELST and DISP contributions, $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$, and the relative difference (in %) between CCSD(T) and DMC. This analysis confirms our prior observations on the trends between DMC and CCSD(T). For example, the more dominant the DISP contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more negative $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to underbind with respect to DMC. Similarly, the stronger the ELST contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more positive $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to overbind with respect to DMC. We expect that this cheap descriptor can be used in the future to identify more challenging systems with larger discrepancies between DMC and CCSD(T).

### v2

@@ Abstract
The accurate treatment of non-covalent interactions is necessary to model a wide range of applications, from molecular crystals to surface catalysts to aqueous solutions and many more. Quantum diffusion Monte Carlo (DMC) and coupled cluster theory with single, double and perturbative triple excitations [CCSD(T)] are considered two widely-trusted methods for treating non-covalent interactions. However, while they have been well-validated for small molecules, recent work has indicated that these two methods can disagree by more than 7.57.57.5\,7.5kcal/mol for larger systems. The origin of this discrepancy remains unknown. Moreover, the lack of systematic comparisons, particularly for medium-sized complexes, has made it difficult to identify which systems may be prone to such disagreements and the potential scale of these differences. In this work, we leverage the latest developments in DMC to compute interaction energies for the entire S66 dataset, containing 66 medium-sized complexes with a balanced representation of dispersion and electrostatic interactions. Comparison to previous CCSD(T) references reveals systematic trends, with DMC predicting stronger binding than CCSD(T) for electrostatic-dominated systems, while the binding becomes weaker for dispersion-dominated systems. We show that the relative strength of this discrepancy is correlated to the ratio of electrostatic and dispersion interactions, as obtained from energy decomposition analysis methods. Finally, we have pinpointed model systems: the hydrogen-bonded acetic acid dimer (ID 20) and dispersion-dominated uracil–cyclopentane dimer (ID 42), where these discrepancies are particularly prominent. These systems offer cost-effective benchmarks to guide future developments in DMC, CCSD(T) as well as the wider electronic structure theory community.
@@ I Introduction
The origin of the discrepancy between DMC and CCSD(T) for large dispersion-bound molecules is a topic of current debate, Schäfer et al. (2024); Fishman et al. (2024); Lambie et al. (2024); Lao (2024) particularly on the validity of the perturbative triples (T) contribution in CCSD(T). Schäfer et al. Schäfer et al. (2024) have suggested that part of this discrepancy arises from missing contributions in (T) that can be accounted by the (cT) approach. In addition, Semidalas et al. Semidalas, Boese, and Martin (2025) have reported non-trivial discrepancies between CCSD(T) and post-CCSD(T) methods such as CCSDT(Q). Conversely, Lambie et al. Lambie et al. (2024) have found that CCSD(T) does not differ significantly against CCSDT(Q) using the Pariser-Parr-Pople (PPP) model Pariser and Parr (1953a, b); Pople (1953) for large conjugated systems. Similarly, Fishman et al. Fishman et al. (2024) and Lao Lao (2024) report only a slight overbinding of CCSD(T) against CCSDT(Q) that cannot explain the discrepancy against DMC.
@@ II.1 Diffusion Monte Carlo
A detailed description of the DMC method can be found in Ref. 9. In this work, we compute fixed-node DMC interaction energies by using the CASINO codeNeeds et al. (2020). We use energy-consistent correlated electron pseudopotentialsTrail and Needs (2017) (eCEPP) with the determinant locality approximation (DLA)Zen et al. (2019). The trial wave-functions were of the Slater–Jastrow type with single Slater determinants, and the single-particle orbitals obtained from DFT local-density approximation (LDA) plane-wave calculations performed with PWscfGiannozzi et al. (2009, 2020) using an energy cut-off of 600 Ry and re-expanded in terms of B-splinesAlfè and Gillan (2004). The Jastrow factor included a two-body electron–electron (e–e) term, two-body electron–nucleus (e–n) terms, and three-body electron–electron–nucleus (e–e–n) terms. The variational parameters of the Jastrow have been optimized by minimizing the variance of each system. The final DMC estimates of $\Delta E_{\text{int.}}$ were extrapolated towards the zero time step limit ($\tau\to 0$) by making a cubic fit to a series of time step estimates from 0.10.10.1\,0.1au down to 0.0030.0030.003\,0.003au. We estimate errors which capture both the stochastic errors in the fit as well as the errors in the cubic fit due to the changing behavior near the zero time step limit. To do this, we make a linear fit on a subset of time steps below (and including) 0.020.020.02\,0.02au and calculate the difference of the extrapolated estimates from the linear fit against the original cubic fit. The final error estimate is taken to be the larger of the stochastic errors of the cubic fit or the difference between the linear fit and cubic fit, as discussed in Sec. S2.4 of the supplementary material.
@@ II.2 Coupled Cluster Theory
Schäfer et al. Schäfer et al. (2024) have recently demonstrated that there exists an empirical relationship for dispersion-dominated complexes between the (cT) and the (T) correlation contributions to the total energy using the CCSD and MP2 correlation energies. The resulting (cT)-fit is of the form:
@@ III Results and Discussion
We report the final DMC estimates for the entire S66 dataset in Figs. 12 and 3. This dataset comprises of a diverse range of interactions and we have separated the systems according to the original S66 categories of hydrogen-bonded, dispersion-bonded and “mixed”-character systems in Figs. 12 and 3, respectively. The corresponding dimer complexes are visualized in Fig. S1 of the supplementary material. We report the DMC estimate of the interaction energy $\Delta E_{\textrm{int.}}$ above the label of each S66 complex. In all cases, the errors on $\Delta E_{\textrm{int.}}$ estimates are below 0.120.120.12\,0.12kcal/mol, with the majority below 0.100.100.10\,0.10kcal/mol, facilitating reliable comparisons to CCSD(T).
@@
The strength of $\Delta E_{\textrm{int.}}$ varies significantly across the systems, from as large as $-20.17\pm 0.07\,$kcal/mol for complex 20 (acetic acid dimer) to as weak as $-1.11\pm 0.06\,$kcal/mol for complex 30 (benzene-ethene dimer), being stronger in the H-bonded systems. With gray crosses, we plot the difference between DMC and CCSD(T) estimates (as described in the Methods) for the three classes of interactions. We use DMC as the reference (i.e., zero), and plot blue vertical bars along the horizontal zero-axis representing the errors on the DMC estimates. There is overall excellent agreement, with a mean absolute deviation (MAD) of 0.210.210.21\,0.21kcal/mol across the entire S66 dataset. We find systematic trends in the differences between CCSD(T) and DMC, with CCSD(T) predicting weaker binding compared to DMC for hydrogen-bonded systems in Fig. 1, with an MAD of $\sim 0.24\,$kcal/mol, while predicting a stronger binding for dispersion dominated systems in Fig. 2, with an MAD of $\sim 0.24\,$kcal/mol. For the “mixed” character systems in Fig. 3, the MAD is lower at 0.140.140.14\,0.14kcal/mol.
Figure 2: Comparison between DMC interaction energies $\Delta E_{\textrm{int.}}$ calculated in the present work against CCSD(T) for a subset of systems in the S66 dataset dominated by dispersion interactions. The first 6 dimers are $\pi{-}\pi$ stacked. Refer to the caption of Fig. 1 for the plot details. The CCSD(T) estimate is taken as an average from three previous calculations Řezáč, Riley, and Hobza (2011b); Kesharwani et al. (2018); Nagy et al. (2023), with corresponding standard deviation as error. Additional CCSD(cT)-fit estimates are reported with golden crosses. These are calculated by scaling the CCSD(T) estimates based on their MP2 and CCSD contributions with the approach described in Ref. 50.
@@
The stronger binding of DMC over CCSD(T) has not been (systematically) reported before, with the acetic acid dimer (ID 20) giving the maximum deviation of 0.80.80.8\,0.8kcal/mol across all S66 systems. Within Sec. S7 of the supplementary material, we have confirmed that the computed $\Delta E_{\textrm{int.}}$ estimate ($-20.17\pm 0.07\,$kcal/mol) does not depend on the chosen pseudopotential (eCEPP), localization scheme (DLA) or trial wave-function (LDA). For example, we have also performed all-electron calculations, giving an estimate of $-20.32\pm 0.12\,$kcal/mol. We have also performed tests using PBE and PBE0 trial wave-functions, showing that the DMC $\Delta E_{\textrm{int.}}$ has a negligible (${<}0.15\,$kcal/mol) dependence on the nodal surface for the DFAs considered. Furthermore, we also computed estimates for two-other localization schemes: T-move and determinant localization T-move, both of which were within the statistical uncertainties of our original estimate. It should be noted that while the absolute value of the difference can be significant for some hydrogen-bonded systems, the relative difference (normalized against the DMC $\Delta E_{\textrm{int.}}$) is significantly smaller, with a mean relative difference of 2.45% compared to 8.21% for the dispersion-dominated systems (see Sec. S6 of the supplementary material).
@@
Figure 4: Error decomposition analysis. We report the difference between DMC and CCSD(T) relative to the DMC magnitude, i.e. $\left(E_{\mathrm{DMC}}-E_{\mathrm{CCSD(T)}}\right)/\left|E_{\mathrm{DMC}}\right|$, as a function of the natural logarithm of the electrostatic (ELST) to dispersion (DISP) ratio contribution to the binding energy. The ELST to DISP ratio is determined from the SAPT analysis from Ref. 94. The color code is red for H-bonded systems (ID from 1 to 23), blue for dispersion dominated systems (ID from 24 to 46), and green for mixed systems (ID from 47 to 66).
@@
The weaker binding of DMC over CCSD(T) for dispersion-dominated systems is now relatively well-documented Al-Hamdani et al. (2021); Schäfer et al. (2024), and there is evidence that it can be improved by replacing the perturbative triples (T) contribution with the recent (cT) Masios et al. (2023) contribution. We plot the difference between an empirically CCSD(cT)-fit formulation as yellow crosses for the dispersion-dominated systems. In all cases, CCSD(cT) has a weaker binding than CCSD(T), leaning closer towards DMC, leading to an MAD of 0.090.090.09\,0.09kcal/mol and mean relative difference of 3.44% However, this does not fully resolve the discrepancies across all of the dispersion systems, with significant discrepancy of ${\sim}0.5\,$kcal/mol remaining for the uracil-cyclopentane dimer (ID 42). Such a significant discrepancy makes this a worthwhile system to investigate further and could give clues on remaining discrepancies between DMC and CCSD(T) observed in Ref. 50. The analysis reported above highlights an important outcome of this work: the identification of smaller, simpler systems that show notable discrepancies between DMC and coupled cluster methods. Specifically, we found a discrepancy of approximately ${\sim}0.8\,$kcal/mol for the hydrogen-bonded acetic acid dimer (ID 20) and ${\sim}0.5\,$kcal/mol for the dispersion-dominated uracil–cyclopentane dimer (ID 42), which contain 64 and 98 (total) electrons, respectively. These medium-sized systems represent an almost tenfold reduction in electron count compared to the larger $\mathrm{C}_{60}@[6]\mathrm{CPPA}$ buckyball-ring system (672 electrons) studied in Ref. 49. Thus, they might offer practical, cost-effective models for further exploring the discrepancy between DMC and coupled-cluster.
@@
We now focus on the difference between DMC and CCSD(T) as a function of the dispersion and electrostatic contribution to the interaction energy. In particular, we find that the relative differences between DMC and CCSD(T) for each system within the S66 dataset can be correlated to the relative strength of the dispersion and electrostatic interactions that make up its $\Delta E_{\textrm{int.}}$. We used the Symmetry Adapted Perturbation Theory (SAPT) calculations (at the $s$SAPT0 level with the jun-cc-pVDZ basis set Parker et al. (2014)) from Burns et al.Burns et al. (2017), which decomposes $\Delta E_{\text{int.}}$ into contributions from electrostatics (ELST), exchange, induction, and dispersion (DISP). Notably, we show in Fig. 4 that there is a strong linear trend ($R^{2}{=}0.78$) between the natural logarithm of the ELST and DISP contributions, $\log\left(\frac{\text{ELST}}{\text{DISP}}\right)$, and the relative difference (in %) between CCSD(T) and DMC. In Sec. S9 of the supplementary material, we show that this strong linear trend remains at the more sophisticated SAPT2+(3)(CCD)/aug-cc-pVTZ level. Villot and Lao (2024) This analysis confirms our prior observations on the trends between DMC and CCSD(T). For example, the more dominant the DISP contribution to $\Delta E_{\text{int.}}$ (i.e., a more negative $\log\left(\frac{\text{ELST}}{\text{DISP}}\right)$), the more CCSD(T) is found to underbind with respect to DMC. Similarly, the stronger the ELST contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more positive $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to overbind with respect to DMC. We expect that this cheap descriptor can be used in the future to identify more challenging systems with larger discrepancies between DMC and CCSD(T).
Finally, we discuss briefly the potential origins of the observed discrepancies between DMC and CCSD(T) based upon the current literature. For H-bonded systems, CCSDT(Q) estimates are available for the A24 dataset Řezáč and Hobza (2013); Řezáč et al. (2015) – a set of $\Delta E_{\text{int.}}$ for small dimer complexes. Nakano et al. Nakano et al. (2024) have performed DMC calculations for the entire A24 dataset, where there is a notable discrepancy of $0.26{\pm}0.07\,$kcal/mol and $0.34{\pm}0.07\,$kcal/mol for the water-ammonia and HCN dimer complexes, in line with our observation of stronger binding in DMC. The CCSDT(Q) references, albeit at small basis sets, find negligible (${<}0.01\,$kcal/mol) changes relative to CCSD(T). For dispersion-bound systems, there exist CCSDT(Q) estimates for the parallel-displaced (PD) benzene dimer by Semidalas et al. Semidalas, Boese, and Martin (2025) and by Karton and Martin Karton and Martin (2021), which report differences to CCSD(T) of $-0.085\,$kcal/mol and $-0.058\,$kcal/mol, respectively, using small truncated double-zeta quality basis sets. The reported difference between DMC and CCSD(T) is $-0.37{\pm}0.08\,$kcal/mol for this system which indicates that the majority of this difference is not covered when going to CCSDT(Q). For both H-bonded and dispersion-bound systems (as well as those of mixed-character), the observed differences could arise either from higher order excitations and larger basis sets needed from coupled cluster theory, or biases in the FN-DMC evaluations, likely coming from the fixed-node approximation.